A company must purchase one reactor to be used in an overall operation. Four reactors have been designed, all of which are equally capable of giving the required service. The following data apply to the four designs:

Fixed-capital (FCI),TL 10000 12000 14000 16000

Sum of operating and fixed costs per year (all other costs are constant) 3000 2800 2350 2100

If the company demands a 15 percent return after taxes on any unnecessary investment, which of the four designs should be accepted? 


1
Expert's answer
2021-05-18T11:21:42-0400
"Solution"

Given the table above:

% of final cost to fixed capital is: "design \ one =30\%"

"design\ two=23.33\%\\
design \ three=16.77\%\\
design \ four=14.13\%"


so design four , housing lowest percentage of cost . it will be selected .


The life of the reactors is not given so we are going to work on infinity term to the perpetuities formula going to be used.

The company should accept the company with the lowest PW of costs .

where :"PW=initial\ cost\ +[\frac{a}{i}]"

"a=Annual \ cost\\
i=interest\ rate"

"PW\ of cost\ of\ design\ one=10000+(\frac{3000}{0.15})=30,000"

"PW\ of\ design\ two=12000+(\frac{ 2800}{0.15})=30,666.6667"


"PW\ of \ design\ three=14000+(\frac{2350}{0.15})=29,666.6667"


"PW \ of \ design \ four=16000+(\frac{2100}{0.15})=30,000"


Design three(3) has the lowest cost . So its should be accepted .